Mary is two years older than Robert. The sum of their ages is 28. Write an expression for each person, write the equation for the situation and find out how old each person is?
I'm confused on where to even start with this question to solve it.

Answer by macston(5194)

R=Robert's age; M=Mary's age=R+2yrs
R+M=28yrs Substitute for M
R+R+2yrs=28yrs Subtract 2 yrs for each side.
2R=26yrs Divide each side by 2.
R=13 yrs ANSWER 1: Robert is 13 years old.
M=R+2yrs=13yrs+2yrs=15yrs ANSWER 2: Mary is 15 years old.